OpenStudy (anonymous):

3/4x+1/4y=13/2 x-3/4y=13/2 solVe linear system~HELP PLEASE~

OpenStudy (anonymous):

3/4x + 1/4y = 13/2. Multiply equation by 4 to eliminate fractions. this gives 3x + y = 26. Next equation x-3/4y = 13/2 -- cross multiplying -- we get 2 (x-3) which is 2x - 6 + (13 x 4y) = 0. Simplifying we get 2x + 52y - 6 = 0. Transposing we get 2x + 52y = 6 (1) ----- x 3 We now have 3x + y = 26 (2) ----- x 2 Multiplying equation (1) by 3 and equation (2) by 2 we get the foll: 6x + 156y = 18 6x + 2y = 52 Eliminating x (by changing the signs in equation 2 and subtracting equation 2 from equation 1) we get 156y -2y = -34. So 154y = -34. y = - 34/154 = -17/87. Substituting value for y in equation 2 we get 3x - 17/87 = 26. 3x = 26 +17/87. x = 26 17/87 /3 = (26 x 87) + 17/87 x 3 = 2262 + 17/261. Ans --- x = 8.73 (y = -17/34)